In April 1974, Steve Prefontaine completed a 10 km race in a time of 27 min, 43.6 s. Suppose Pre was at the 8.27 km mark at a time of 25.0 min. If he accelerated for 60 s and then maintained his increased speed for the remainder of the race, calculate his acceleration over the 60 s interval. Assume his instantaneous speed at the 8.27 km mark was the same as his overall average speed up to that time. acceleration: 0.181 m/s2 contact us help 8
Answer #1

Speed =Distance/time

Initial speed (Speed at 8.27km mark) =8.27km/25min

u=0.3308km/min= 5.513m/s

Accelerated for 60 s to achieve increased speed

v=u+at

v=5.513m/s+a*60s

{\color{Red} v=5.513+60a} -----------------(1)

-------------------------------------

Distance covered while accelerating

s=ut+1/2at^{2}

s=5.513m/s*60s+1/2a(60s)^{2}

{\color{Red} s=330.78+1800a} -----------(2)

----------------------------

Time remaining = 27min 43.6s - 25min -60s =103.6s
Remaining distance =10km -8.27km - s=d

1730m-s=d

Remaining distance is covered  using increased speed

v=d/103.6

v=(1730-s)/103.6

From (1)   5.513+60a =(1730-s)/103.6

(5.513+60a) 103.6=1730-s

s=1730-(5.513+60a) 103.6 -------------(3)

Put (3) in (2)

1730-(5.513+60a) 103.6=330.78+1800a

1730-571.147-6216a=330.78+1800a

1730-571.147-330.78=1800a+6216a

828.073=8016a

ANSWER: {\color{Red} a=0.1033m/s^{2}}

=========================